# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 1.3. Outline of the paper
The intrinsic mirror family to a quasi-projective open Kulikov degeneration $\mathcal{X}\to\Delta$ naturally extends to a family $\mathcal{Y}_{\mathcal{X}}\to S_{\mathcal{X}}$ over the affine toric variety $S_{\mathcal{X}}:=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$, with fan the nef cone $\mathrm{Nef}(\mathcal{X}/\overline{\mathcal{X}})$.
The proof of this uses the birational geometry of the crepant resolution $\mathcal{X}\to\overline{\mathcal{X}}$ to constrain the curve classes appearing in the intrinsic mirror construction.
By construction [48], the intrinsic mirror algebra $\mathcal{R}_{\mathcal{X}}$ admits a basis of theta functions $\{\vartheta_{p}\}_{p\in C(P)_{\mathbb{Z}}}$, indexed by the integral points of the tropicalization of $\mathcal{X}$ given by the cone $C(P)$ over $P$. It follows by Lemma 6.22 that the theta functions $\vartheta_{p}$ corresponding to the interior integral points of $C(P)$ define an ideal of the mirror algebra $R_{\mathcal{X}}$. We prove that this ideal defines an anticanonical divisor $\mathcal{D}_{\mathcal{X}}$ in $\mathcal{Y}_{\mathcal{X}}$. On the other hand, the intrinsic mirror family is naturally polarized by the line bundle $\mathcal{L}_{\mathcal{X}}=\mathcal{O}_{\mathcal{Y}_{\mathcal{X}}}(1)$ with a natural basis of sections $\vartheta_{p}$, for $p\in P_{\mathbb{Z}}$, and so we have a natural “theta divisor” $\mathcal{C}_{\mathcal{X}}:=\{\sum_{p\in P_{\mathbb{Z}}}\vartheta_{p}=0\}\in|\mathcal{L}_{\mathcal{X}}|$. Hence, running the intrinsic mirror construction for every projective crepant resolution $\mathcal{X}^{\prime}\rightarrow\overline{\mathcal{X}}$ of the affinization $\overline{\mathcal{X}}$ of $\mathcal{X}$, we obtain pairs $\left(\mathcal{Y}_{\mathcal{X}^{\prime}},D_{\mathcal{X}^{\prime}}+\epsilon\,\mathcal{C}_{\mathcal{X}^{\prime}}\right)$ over the affine toric varieties $\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}^{\prime}/\overline{\mathcal{X}})]$. In §6 we show that these families glue to a family over the movable secondary fan and in Theorem 6.28 we show that this family extends uniquely to a family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\to\mathcal{S}^{sec}_{\mathcal{X}}$, where $\mathcal{S}^{sec}_{\mathcal{X}}$ is a toric Deligne-Mumford stack defined by a stacky fan $\mathcal{S}ec(\mathcal{X}/\overline{\mathcal{X}})$ with underlying fan the secondary fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$. Theorems 7.13 and 8.10 then prove:
@@
To prove that the fibers $(Y_{t},D_{t}+\epsilon\,C_{t})$ of the intrinsic mirror family are slc, we first show in Theorem 7.7 that over the dense torus of the toric stack $\mathcal{S}_{\mathcal{X}}^{\mathrm{sec}}$, the surface $Y_{t}$ is irreducible, and then it is enough to show that $C_{t}$ does not pass through the 00-strata of $D_{t}$. This follows from the description in §6.4.2 of the restriction of the theta functions to $D_{t}$. For an arbitrary fiber $(Y_{t},D_{t}+\epsilon\,C_{t})$, with irreducible components $(V_{i},D_{i}+\epsilon\,C_{i})$, we show using Theorem 7.12 that it suffices to prove that the pairs $(V_{i},D_{i}+\epsilon\,C_{i})$ are slc. By the naturalness of the intrinsic mirror symmetry, this follows from Theorem 7.7.
To study the general fiber of the intrinsic mirror family to the semistable mirror $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$, we show that, up to a base change of $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$, a particular one-parameter sub-family of the intrinsic mirror family is a toric degeneration in the sense of Gross–Siebert [47]. This allows us to identify the general fiber topologically, and then as a log Calabi–Yau surface using a result of Friedman [34]. Then, by the universal property of the moduli space $\mathcal{M}_{(Y,D,L)}$, we obtain a map of stacks $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\rightarrow\mathcal{M}_{(Y,D,L)}$, inducing a map $S^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\rightarrow M_{(Y,D,L)}$ at the level of coarse moduli spaces. To conclude our main result, in Theorem 8.11 we show that this map is finite and surjective. The finiteness follows by proving that the intrinsic mirror family is non-constant in restriction to the one-dimensional strata of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$, which suffices by [50, Lemma 8.7]. Finally, the surjectivity is proved using a period calculation, as in [71, 83].
@@ 1.4. Related work
Mirrors to log smooth maximal degenerations to (log) Calabi–Yau varieties are constructed by Gross–Siebert in [48]. As the maximal degenerations of polarized log Calabi–Yau surfaces we consider in this paper are typically not log smooth, the construction of [48] does not apply directly to produce the semistable mirror $\mathcal{X}\rightarrow\Delta$. Also, the mirror construction of Gross–Hacking–Keel [41] concerns a single “smooth” log Calabi–Yau surface, without maximal degenerations, and hence it does not apply either. To address this issue, we give a direct construction by deformation theory of the semistable mirror $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$. If we restrict attention to smooth log Calabi–Yau surfaces, we expect the affinization $\mathcal{\overline{X}}_{\mathscr{P}}$ of our semistable mirror $\mathcal{X}_{\mathscr{P}}$ to be deformation equivalent to the mirror of [41]. However, the existence of a crepant resolution in that setup is not immediate. Moreover, in order to eventually relate the geometry of the “double mirror” to the original polarized log Calabi–Yau surface, the semistable mirror construction introduced in our paper plays a crucial role.
@@
To construct the double mirror to $(Y,D,L)$, that is, the intrinsic mirror to $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$, we use a version of the Gross–Siebert mirror construction [48, 49], which is possible since $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ is log smooth. In [50], the non-Archimedean mirror construction developed by Keel–Yu [64] is used instead to verify the Hacking–Keel–Yu conjecture for $6$ stable log Calabi–Yau surfaces $(Y,D+\epsilon\,C)$, namely when $Y$ is a del Pezzo surfaces of degree $n$, for $1\leq n\leq 6$, endowed with the line bundle $L=-K_{Y}$, and $D$ a cycle of $n$ $(-1)$-curves. These surfaces are in a sense “self-mirror”, for example, the dimension of the complex moduli equals the dimension of the Kähler moduli, and due to this coincidence one can conveniently take the mirror to any of these del Pezzo’s $Y$ as the canonical bundle $K_{Y}$. In this paper, to prove the Hacking–Keel–Yu conjecture for all log Calabi–Yau surfaces, we investigate double mirrors by studying intrinsic mirrors to our semistable mirrors, applying (a version of) the construction of [48] rather than the non-Archimedean mirror construction of [64]. To be able to use the non-Archimedean mirror construction in this generality, it would first be necessary to have an extension of the results of [64] over a field with nontrivial valuation, which is not yet available [^1] After the second author presented results of this paper at the Homological Mirror Symmetry conference on January 28th 2024 at IMSA/Miami, we were informed that Logan White is currently working on such an extension in his PhD under the guidance of Sean Keel..
Finally, the ongoing work of Gross–Hacking–Keel–Siebert [44] on describing compactifications of moduli spaces of K3 surfaces using intrinsic mirror symmetry is also closely related, as open Kulikov degenerations introduced in our paper are non-compact analogues of the Type III Kulikov degenerations of K3 surfaces – see also [54] and [55] in this context.
@@ 1.5. Acknowledgments
Valery Alexeev was partially supported by the NSF grant DMS-2201222. Hülya Argüz was partially supported by the NSF grant DMS-2302116, and Pierrick Bousseau was partially supported by the NSF grant DMS-2302117. We thank the organizers of the “Homological Mirror Symmetry” and the “Geometry of the South” conferences at IMSA/Miami, where this work was presented, for their hospitality. Finally, we thank Philip Engel for many useful discussions at an early stage of this project.
@@ 2. Polarized log Calabi–Yau surfaces and Symington polytopes
In this section, we first review in §2.1 the definition and basic properties of polarized log Calabi–Yau surfaces. Then, we explain in §2.2, following [31], how to associate to such surfaces “nice” choices of “Symington polytopes”. The main result of this section is the existence of “good” polyhedral decompositions on such polytopes, obtained in Theorem 2.14 in §2.3. Finally, we describe in §2.4 how to construct maximal degenerations of polarized log Calabi–Yau surfaces from good polyhedral decompositions on Symington polytopes.
@@ 2.1. Log Calabi–Yau surfaces
In this paper, a log Calabi–Yau surface is a log canonical (lc) pair $(Y,D)$ over $\mathbb{C}$, as in [67, Definition 2.34], with $Y$ projective of dimension $2$, $D$ a reduced effective divisor, and $K_{Y}+D$ trivial, in particular Cartier. Throughout this paper we assume that all log Calabi–Yau surfaces are maximal, in the sense that $D$ is non-empty and singular. The following proposition classifies singularities of log Calabi–Yau surfaces.
@@
    $(\mathbb{A}^{2}_{x,y}/\frac{1}{r}(1,a),(xy=0))$, where $r$ and $a$ are coprime integers, with $r\geq 2$ and $11\leq a\leq r-$ 1: $Y$ has a cyclic quotient singularity at $p$, obtained as the quotient of $\mathbb{A}^{2}_{x,y}$ by $\mathbb{Z}/r\mathbb{Z}$ acting by $x\to\zeta x$, $y\to\zeta^{a}y$, where $\zeta$ is a primitive r’th root of $1$. Moreover, locally around $p$, $D$ is the image under the quotient by $\mathbb{Z}/r\mathbb{Z}$ of the union of the coordinate axes in $\mathbb{A}^{2}_{x,y}$, and consequently $D$ has an ordinary double point at $p$.
@@
See [50, Propositions 7.2–7.3]. ∎
Note that while $(Y,D)$ is in general singular, we can often work with its minimal resolution.
@@
The statement follows from Proposition 2.1 and the toric description of minimal resolutions of cyclic quotient singularities [25, Chapter 10]. ∎
@@
We will consider polarized log Calabi–Yau surfaces $(Y,D,L)$, given by a log Calabi–Yau surface $(Y,D)$ endowed with an ample line bundle $L$. In what follows we describe toric models for such surfaces. For this, we first define first define interior and corner blow-ups.
@@
Every Calabi-Yau surface admits a toric model by [41, Lemma 1.3] applied to the minimal resolution given by Corollary 2.2. We will use particularly good toric models, defined as follows.
@@
Every polarized log Calabi–Yau surface $(Y,D,L)$ admits a good toric model.
@@
This follows from the proof of [31, Theorem 5.4] applied to the smooth log Calabi–Yau pair $(Y^{\prime},D^{\prime})$ obtained as the minimal resolution $g^{\prime}:(Y^{\prime},D^{\prime})\rightarrow(Y,D)$ as in Corollary 2.2, and to the big and nef line bundle $L^{\prime}:=(g^{\prime})^{\star}L$. Note that, while [31, Theorem 5.4] is stated for an ample line bundle, its proof applies more generally to a big and nef line bundle as remarked in [31, Remark 5.7]. ∎
@@
By the existence of a toric model, it follows that $Y$ is isomorphic to a projective toric surface $\overline{Y}$, up to loci of complex dimension one, that is, of real dimension two. Hence, we have $H_{1}(Y,\mathbb{Z})=H_{1}(\overline{Y},\mathbb{Z})=0$, so $H^{2}(X,\mathbb{Z})$ is torsion free, and $\mathrm{Pic}(Y)$ is torsion free by the exponential exact sequence. ∎

### v3

@@ 1.3. Outline of the paper
The intrinsic mirror family to a quasi-projective open Kulikov degeneration $\mathcal{X}\to\Delta$ naturally extends to a family $\mathcal{Y}_{\mathcal{X}}\to S_{\mathcal{X}}$ over the affine toric variety $S_{\mathcal{X}}:=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]$, with fan the nef cone $\mathrm{Nef}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$.
The proof of this uses the birational geometry of the crepant resolution $\mathcal{X}\to\mathcal{X}^{\mathrm{can}}$ to constrain the curve classes appearing in the intrinsic mirror construction.
By construction, the intrinsic mirror algebra $\mathcal{R}_{\mathcal{X}}$ admits a basis of theta functions $\{\vartheta_{p}\}_{p\in C(P)_{\mathbb{Z}}}$, indexed by the integral points of the tropicalization of $\mathcal{X}$ given by the cone $C(P)$ over $P$. Lemma 6.33 establishes that the theta functions $\vartheta_{p}$ corresponding to the interior integral points of $C(P)$ generate an ideal of the mirror algebra $R_{\mathcal{X}}$. Following [53, §6.2], we prove that this ideal defines an anticanonical divisor $\mathcal{D}_{\mathcal{X}}$ in $\mathcal{Y}_{\mathcal{X}}$. On the other hand, the intrinsic mirror family is naturally polarized by the line bundle $\mathcal{L}_{\mathcal{X}}=\mathcal{O}_{\mathcal{Y}_{\mathcal{X}}}(1)$ with a natural basis of sections $\vartheta_{p}$, for $p\in P_{\mathbb{Z}}$. Consequently, as in [53, §6.2], we have a natural “theta divisor” $\mathcal{C}_{\mathcal{X}}:=\{\sum_{p\in P_{\mathbb{Z}}}\vartheta_{p}=0\}\in|\mathcal{L}_{\mathcal{X}}|$. Hence, applying the intrinsic mirror construction for every projective crepant resolution $\mathcal{X}^{\prime}\rightarrow\mathcal{X}^{\mathrm{can}}$ of the affinization $\mathcal{X}^{\mathrm{can}}$ of $\mathcal{X}$, we obtain pairs $\left(\mathcal{Y}_{\mathcal{X}^{\prime}},D_{\mathcal{X}^{\prime}}+\epsilon\,\mathcal{C}_{\mathcal{X}^{\prime}}\right)$ over the affine toric varieties $\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})]$. In §6 we build upon arguments of [53, §6] to show that these families glue together into a larger family over the movable secondary fan. Furthermore, in Theorem 6.41, we establish that this family extends uniquely to a family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\to\mathcal{S}^{sec}_{\mathcal{X}}$, where $\mathcal{S}^{sec}_{\mathcal{X}}$ is a toric Deligne-Mumford stack defined by a stacky fan $\mathcal{S}ec(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, whose underlying fan is the secondary fan $\mathrm{Sec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. Theorems 7.7 and 8.11 then prove:
@@
To prove that the fibers $(Y_{t},D_{t}+\epsilon\,C_{t})$ of the intrinsic mirror family are semi-log-canonical, we first consider the case of fibers over the dense torus of the toric stack $\mathcal{S}_{\mathcal{X}}^{\mathrm{sec}}$. In this case, we first show in Theorem 7.1 that, as in [53, §7], the surface $Y_{t}$ is irreducible. Then, to prove that $(Y_{t},D_{t}+\epsilon\,C_{t})$ is semi-log-canonical, it suffices to verify that $C_{t}$ does not intersect the 0-strata of $D_{t}$. This follows from the description in §6.4.4 of the restriction of the theta functions to $D_{t}$. For an arbitrary fiber $(Y_{t},D_{t}+\epsilon\,C_{t})$, with irreducible components $(V_{i},D_{i}+\epsilon\,C_{i})$, we show using Theorem 7.6 that proving that the pairs $(V_{i},D_{i}+\epsilon\,C_{i})$ are semi-log-canonical is sufficient. By the naturalness of the intrinsic mirror symmetry, this follows from Theorem 7.1.
To study the general fiber of the intrinsic mirror family to the semistable mirror $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$, we show that, up to a base change of $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$, a specific one-parameter sub-family of the intrinsic mirror family constitutes a toric degeneration in the sense of Gross–Siebert [50]. This allows us to first determine the topology of the general fiber, and then to identify it as a log Calabi–Yau surface using a result of Friedman [36]. Then, by the universal property of the moduli space $\mathcal{M}_{(Y,D,L)}$, we obtain a map of stacks $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\rightarrow\mathcal{M}_{(Y,D,L)}$, which induces a corresponding map $S^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\rightarrow M_{(Y,D,L)}$ at the level of coarse moduli spaces. To conclude our main result, Theorem 8.13 demonstrates that this map is finite and surjective. The finiteness follows from [53, Lemma 8.7] by proving that the intrinsic mirror family is non-constant in restriction to the one-dimensional strata of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$. Finally, the surjectivity is established through a dimension calculation.
@@ 1.4. Related work
Mirrors to log smooth maximal degenerations to (log) Calabi–Yau varieties are constructed by Gross–Siebert in [51]. As the maximal degenerations of polarized log Calabi–Yau surfaces we consider in this paper are typically not log smooth, the construction of [51] does not apply directly to produce the semistable mirror $\mathcal{X}\rightarrow\Delta$. Also, the mirror construction of Gross–Hacking–Keel [44] concerns a single “smooth” log Calabi–Yau surface, without maximal degenerations, and hence it does not apply either. To address this issue, we give a direct construction by deformation theory of the semistable mirror $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$. If we restrict attention to smooth log Calabi–Yau surfaces, we expect the affinization $\mathcal{X}^{\mathrm{can}}_{\mathscr{P}}$ of our semistable mirror $\mathcal{X}_{\mathscr{P}}$ to be deformation equivalent to the mirror of [44] – see Remark 4.8.
@@
To construct the double mirror to $(Y,D,L)$, that is, the intrinsic mirror to $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$, we use a version of the Gross–Siebert mirror construction [51, 52], which is possible since $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ is log smooth. In [53, 54], the non-Archimedean mirror construction developed by Keel–Yu [70] is used in arbitrary dimension to construct intrinsic mirror families over the toric variety defined by the secondary fan of crepant resolutions $\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$, which are algebraic over $\Delta=\mathbb{A}^{1}$. In the present paper, we use instead in dimension two the Gross–Siebert mirror construction, which can be used for crepant resolutions over the formal disk $\Delta=\mathrm{Spf}\,\mathbb{C}[\![t]\!]$. To be able to use the non-Archimedean mirror construction in this set-up, it would first be necessary to have an extension of the results of [70] for formal families. [^1] After the second author presented results of this paper at the Homological Mirror Symmetry conference on January 28th 2024 at IMSA/Miami, we were informed that Logan White is currently working on such an extension in his PhD under the guidance of Sean Keel.
In [53, Theorem 1.3], the general construction of the non-Archimedean intrinsic mirror family over the toric variety defined by the secondary fan is used to prove the full Hacking–Keel–Yu conjecture on $\mathcal{M}_{(Y,D,L)}$ for the six stable log Calabi–Yau surfaces $(Y,D+\epsilon\,C)$, where $Y$ is a del Pezzo surfaces of degree $n$, for $1\leq n\leq 6$, endowed with the line bundle $L=-K_{Y}$, and $D$ a cycle of $n$ (1)(-1)-curves. These surfaces are in a sense “self-mirror”, for example, the dimension of the complex moduli equals the dimension of the Kähler moduli, and due to this coincidence one can conveniently take the mirror to any of these del Pezzo’s $Y$ as the canonical bundle $K_{Y}$. In this paper, to prove the Hacking–Keel–Yu conjecture for all log Calabi–Yau surfaces, a key result is the construction of the semistable mirror $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$, which can then be taken as an input of a general intrinsic mirror construction. Moreover, the semistable mirror is explicit enough to allow us to prove that the “double mirror” is related to the original $(Y,D,L)$. The ongoing work of Gross–Hacking–Keel–Siebert [47] on describing compactifications of moduli spaces of K3 surfaces using intrinsic mirror symmetry is also closely related, as open Kulikov degenerations introduced in our paper are non-compact analogues of the Type III Kulikov degenerations of K3 surfaces – see also [59] and [60] in this context.
In [6], we apply the semistable mirror construction of the present paper to answer a long-standing question in the string theory literature on canonical 3-fold singularities which are M-theory dual to the 5d superconformal field theories constructed from webs of 5-branes with 7-branes. Finally, we will show in future work that the toric variety $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is independent of the choice of the Symington polytope $(P,\mathscr{P})$. We will also give a modular interpretation of the finite morphism in Theorem A.
@@ 1.5. Acknowledgments
Valery Alexeev was partially supported by the NSF grant DMS-2201222. Hülya Argüz was partially supported by the NSF grant DMS-2302116, and Pierrick Bousseau was partially supported by the NSF grant DMS-2302117. We thank the organizers of the “Homological Mirror Symmetry 2024” and the “Geometry of the South” conferences at IMSA/Miami, where this work was presented, for their hospitality. We also thank Philip Engel for many useful discussions at an early stage of this project. Finally, we thank the anonymous referees for their careful reading and the many suggestions to improve the exposition.
@@ 2. Polarized log Calabi–Yau surfaces and Symington polytopes
In this section, we first review the definition and basic properties of polarized log Calabi–Yau surfaces in §2.1. Next, following [33], we explain how to associate to such surfaces “nice” choices of “Symington polytopes” in §2.2. The main result of this section is the existence of “good” polyhedral decompositions on such polytopes, obtained in Theorem 2.16 in §2.3. Finally, in §2.4, we outline the construction of maximal degenerations of polarized log Calabi–Yau surfaces from good polyhedral decompositions on Symington polytopes.
@@ 2.1. Log Calabi–Yau surfaces
In this paper, a log Calabi–Yau surface is a log canonical pair $(Y,D)$ over $\mathbb{C}$, as in [73, Definition 2.34], with $Y$ projective of dimension $2$, $D$ a reduced effective divisor, and $K_{Y}+D$ trivial, in particular Cartier. Throughout this paper we assume that all log Calabi–Yau surfaces are maximal, in the sense that $D$ is non-empty and singular. The following proposition classifies singularities of log Calabi–Yau surfaces.
@@
    $(\mathbb{A}^{2}_{x,y}/\frac{1}{r}(1,a),(xy=0))$, where $r$ and $a$ are coprime integers, with $r\geq 2$ and $1\leq a\leq r-1$: $Y$ has a cyclic quotient singularity at $p$, obtained as the quotient of $\mathbb{A}^{2}_{x,y}$ by $\mathbb{Z}/r\mathbb{Z}$ acting by $x\to\zeta x$, $y\to\zeta^{a}y$, where $\zeta$ is a primitive r’th root of $1$. Moreover, locally around $p$, $D$ is the image under the quotient by $\mathbb{Z}/r\mathbb{Z}$ of the union of the coordinate axes in $\mathbb{A}^{2}_{x,y}$, and consequently $D$ has an ordinary double point at $p$.
@@
See [53, Propositions 7.2–7.3]. ∎
Although a log Calabi–Yau surface $(Y,D)$ is in general singular, we can often work with its minimal resolution.
@@
The statement follows from Proposition 2.1 and the toric description of minimal resolutions of cyclic quotient singularities [26, Chapter 10]. ∎
@@
We will consider polarized log Calabi–Yau surfaces $(Y,D,L)$, given by a log Calabi–Yau surface $(Y,D)$ endowed with an ample line bundle $L$. In what follows we describe toric models for such surfaces. For this, we first define interior and corner blow-ups.
@@
Every Calabi–Yau surface admits a toric model by [44, Lemma 1.3] applied to the minimal resolution given by Corollary 2.2.
As in [45, Definition 1.4], a log Calabi–Yau surface $(Y,D)$ is called generic if its minimal resolution $(Y^{\prime},D^{\prime})$ does not contain an internal (2)(-2)-curve, that is a smooth rational curve $C$ with $C^{2}=-2$ and disjoint from $D^{\prime}$. By [45, Proposition 4.1], every log Calabi–Yau surface is deformation equivalent to a generic one. For generic log Calabi–Yau surfaces, we will use particularly good toric models, defined as follows.
@@
Every generic polarized log Calabi–Yau surface $(Y,D,L)$ admits a good toric model.
@@
This follows from the proof of [33, Theorem 5.4] applied to the smooth log Calabi–Yau pair $(Y^{\prime},D^{\prime})$ obtained as the minimal resolution $g^{\prime}:(Y^{\prime},D^{\prime})\rightarrow(Y,D)$ as in Corollary 2.2, and to the big and nef line bundle $L^{\prime}:=(g^{\prime})^{\star}L$. Note that, while [33, Theorem 5.4] is stated for an ample line bundle, its proof applies more generally to a big and nef line bundle as remarked in [33, Remark 5.7]. ∎
@@
By the existence of a toric model, $Y$ is isomorphic to a projective toric surface $\overline{Y}$, up to loci of complex dimension one, that is, of real dimension two. Since proper toric varieties have trivial first homology groups, we obtain $H_{1}(Y,\mathbb{Z})=H_{1}(\overline{Y},\mathbb{Z})=0$. It follows that $H^{2}(X,\mathbb{Z})$ is torsion free, and by the exponential exact sequence, $\mathrm{Pic}(Y)$ is also torsion free. ∎